Problem 10-10 Calculating Project NPV [LO1]

Quad Enterprises is considering a new three-year expansion project that requires an initial fixed asset investment of $2.29 million. The fixed asset will be depreciated straight-line to zero over its three-year tax life, after which time it will be worthless. The project is estimated to generate $1,715,000 in annual sales, with costs of $625,000. The tax rate is 21 percent and the required return on the project is 10 percent. What is the project’s NPV? (Do not round intermediate calculations. Enter your answer in dollars, not millions of dollars, and round your answer to 2 decimal places, e.g., 1,234,567.89.)
Answer #1

Annual depreciation=(Cost-Salvage value)/Useful Life

=(2,290,000/3)=$763333.333

OCF=(Sales-Costs)(1-tax rate)+Tax savings on Annual depreciation

=(1,715,000-625,000)(1-0.21)+(0.21*763333.333)

=$1021400

Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=1021400[1-(1.1)^-3]/0.1

=1021400*2.48685199

=$2540070.62

NPV=Present value of inflows-Present value of outflows

=2540070.62-2,290,000

=$250070.62(Approx).